# Not flexible enough? Impacts of electric carsharing on a power sector with variable renewables11footnotemark: 1

### v1

@@ 2.3 Power system modelling
[…]
Modeling interactions with BEVs
Sector coupling with BEVs is modelled using distinct vehicle profiles. These profiles represent individual vehicles which are scaled up to the overall fleet. In the case of uncontrolled charging, each profile is considered an exogenous electricity demand that is added to the overall demand. When considering smart or bidirectional charging, the model chooses when and how much electricity should flow into a vehicle, while ensuring battery levels that are compatible with the mobility constraint. This relies on the exogenous time series “driving electricity consumption” and “grid availability” derived with emobpy. Discharging electricity from BEV back to the grid incurs additional variable costs of 5 euros/MWh to account for battery degradation.
The model is calibrated based on a 15 million BEV fleet in 2030, which aligns with the target of the German coalition. The distribution of cars across locations is assumed to be the same as during the survey (Table 6). Additionally, it is assumed that clusters with very long trips (clusters 1 and 2 in each location) are not regular daily patterns, but rather trips undertaken under exceptional circumstances, such as holidays. As these patterns represent a very small proportion of overall sequences (around 2.4%), we exclude them from our analysis and assume that the 15 million BEV fleet is distributed across locations and clusters 3 to 6 (Table 1).
@@ 2.4 Scenarios
We investigate a number of scenarios which differ with respect to the size of the overall vehicle fleet, the number of shared cars, and the charging strategy. Scenarios are compared to respective reference settings without carsharing, i.e., references only include privately-owned BEV profiles.
Scenario families
We group scenarios into two broad families. The first family only considers BEVs that switch to carsharing in the scenario (“Shared only”). The second family additionally includes other electric cars that stay private both in the reference and the scenario (“Shared + other BEVs”). Hence, the BEV fleets considered in the latter family are way larger than the ones in the former (Table 2). The former family highlights the power sector effects of switching to carsharing, assuming away many other potential sources sources of flexibility. While results are instructive, they may not be policy-relevant. The latter family is designed to mimic a more realistic setting with additional BEVs that do not switch to car sharing, but remain privately-owned. Depending on the charging strategy, these additional “non-switching” BEVs also serve as a proxy for other flexibility options that are not included in the model. We thus consider “Shared + other BEVs” to be the main scenario family in the following.
@@
Table 2: BEV units, battery capacity and power consumption for all reference settings and scenarios
@@
For each scenario family, we consider two possible carsharing uptake regimes: a low and a high one. In the low uptake regime, only BEVs belonging to the cluster of shortest trips (cluster 6) are assumed to switch to carsharing. In the high uptake regime, we assume that all BEVs in metropolises switch to carsharing as well as most BEVs in big cities and some in the middle-size cities. In small cities, only cluster 6 is assumed to switch to carsharing. For both uptake regimes, we assume that there is no switch to carsharing in rural areas. The number of substituted private BEVs is approximately 3.9 million in the low uptake scenarios compared to 6.2 million in the high uptake scenarios. Further details are available in Figure 2 and Table 2. The low uptake regime is more conservative, as only users that rely the less on their cars switch to carsharing. The high uptake regime is more progressive and aims at illustrating a very optimistic scenario when analyzing impacts on the power system. Average grid availability for each hour of the week is illustrated in Figure 9.
Figure 2: Definition of carsharing uptake regimes based on cluster and location types. Note that this represents the reference fleet of the “Shared + other BEVs” scenario family. For the “Shared only” family, the reference fleet is only constituted of the shaded areas that switch to shared BEVs in the scenario.
@@
For each scenario, we consider various charging strategies: uncontrolled charging, smart charging, or bidirectional charging. To not mix up effects, we assume that both the reference and the respective scenarios always have the same charging strategy. Additionally, only one charging strategy is used for the entire fleet at a time to enable meaningful comparisons. This means that all BEVs charge smartly in the reference and the scenarios, or they all charge bidirectionally, etc. We consider different charging strategies since it is unclear which charging strategy will be mostly used by BEVs in the future. While it is unlikely that the whole BEV fleet will be operated in an uncontrolled manner (schill2015power (36)), it also seems unlikely that the whole fleet will be operated bidirectionally. By considering these strategies and homogeneous fleets, we analyze extreme settings which can be thought of as boundaries for outcomes of alternative scenarios with BEV fleets with mixed charging strategies.
@@ 3 Results
In the main part of the paper, we focus on the main scenario family where only part of the total BEV fleet considered in the reference switches to carsharing (“Shared + other BEVs”), i.e., the more policy-relevant scenario family. Results for the “Shared only” scenario family are provided in the appendix (D). Qualitative results are similar across scenario families unless explicitly mentioned.
@@ 3.1 Power sector cost increases remain moderate and are largest for bidirectional charging scenario
The adoption of carsharing in the BEV fleet results in increased power system costs across almost all scenario families, carsharing uptake regimes, and charging strategies (Figure 3 and Table 9). This cost increase is relatively small compared to the overall sector costs, and is always below 100 euros per substituted car per year in the main scenario family “Shared + other BEVs”, even in the case with bidirectional charging. In the “Shared + other BEVs” framework, it represents about 1.1% of the total sector costs and 2.2% in the “Shared only” framework. Note that the load of substituted BEVs is also small compared to the overall load and ranges between 1.1% and 2.6% of the overall load in the “Shared + other BEVs” scenario family, depending on the uptake regime.
@@
Figure 3: Power sector costs by charging strategy and carsharing uptake (“Shared + other BEVs”)
@@
In the case where all BEVs charge in an uncontrolled way, switching to electric carsharing implies very little additional cost or may even lead to slight cost reductions. This is because charging a large fleet of BEVs without considering power sector dynamics leads to a very expensive power sector, as BEVs considerably increase existing system peak loads (Figures 12 and 13). By switching to carsharing, the changing charging patterns may even slightly ease this burden. The fact that cost increases related to carsharing remain small should not be misinterpreted. The absolute sector costs remain the highest compared to other charging strategies, even after a switch to carsharing. Since the perspective of having a large BEV fleet charged in an uncontrolled way is neither realistic nor desirable (compare schill2015power (36)), we do not focus on this charging strategy in the rest of the paper. Results can nevertheless be found in the appendix (D).
@@
Assuming that the whole BEV fleet charges smartly or bidirectionally, costs in the reference decrease compared to a setting with uncontrolled charging, as the former charging strategies help avoid costly load peaks and corresponding peaker plants and storage capacity (see section 3.2). In turn, switching to carsharing then increases costs, regardless of the uptake regime. This is because substituting private BEVs that are charged smartly or bidirectionally with shared BEVs partly takes away the flexibility that private BEVs provide to the power sector, as shared cars are driven more often. Note that we assume that private cars provide flexibility to the power sector with no additional investment costs. Substituting these private cars away implies an increase in the investment needed in the power sector. The cost increase is about twice as high when the whole BEV fleet is charged bidirectionally compared to when it is charged smartly, independent of the modeled uptake regime.
@@
For both the smart and the bidirectional charging strategy, the increase in cost is higher when the carsharing uptake is higher. In the “Shared + other BEVs” scenario family, this holds not only in absolute terms, but also in relative terms: the cost increase in euros per substituted car is always greater, everything else being equal, when considering the high uptake. It ranges between 42 and 54 euros per substituted BEV in the smart charging case, and between 92 and 98 euros for bidirectional charging.
@@
The greater the available flexibility options, the lower the cost increase due to a switch to carsharing. For instance, the increase in power sector cost is higher across all charging strategies and uptake rates when considering the “Shared only” framework, where no private BEV remain as a potential flexibility resource (Table 9). Private BEVs (which are absent in the “Shared only” scenario family) are more likely to be used flexibly since they are driven less and are thus available for grid interactions more often. The contrast between the two families is particularly clear-cut for the bidirectional charging strategy, where the cost increase per substituted car more than doubles in “Shared only” compared to “Shared + other BEVs”. The respective increase is only about 1.2 when assuming that all BEVs only charge smartly. This is because feeding back electricity into the grid is particularly useful in power sectors with high shares of renewables, as it helps to substitute peak generation and electricity storage capacities. Reducing the potential of a BEV fleet to engage in V2G accordingly hurts more than just reducing its potential to charge smartly.
@@
Figure 4: Generation and storage discharging capacity by technology (“Shared + other BEVs”)
@@ 3.2 Changes in generation capacity and dispatch depend on charging strategy
The switch to carsharing triggers changes in the optimal electricity generation capacity mix that are different for smart and bidirectional charging, while the minimum renewable share constraint of 80% is binding. In the case of smart charging, the switch to carsharing reduces the potential to flexibilize the demand. In the case of bidirectional charging, switching to carsharing also entails a loss in the ability of the BEV fleet to feed back electricity into the grid when the demand is relatively high compared to the available supply. In all cases, the optimal PV generation capacity increases, by between around 2 and 7 GW (Figure 4, see also Figure 10).
@@
In contrast, optimal wind power capacities generally decrease with carsharing. This is because the relatively large cumulative battery capacity of the fleet of privately owned BEV fosters the integration of wind power in the reference. This is visible in the cumulative BEV battery charging level time series, which not only shows diurnal patterns related to PV variability, but is dominated by longer-duration patterns that relate to wind power variability (Figures 24 and 25). If part of the BEV fleet switches to carsharing, the potential for such longer-duration storage decreases in line with the shrinking overall battery capacity. In turn, the share of wind power in the optimal capacity mix decreases, while PV with its shorter-duration diurnal variations increases. The only exception is the low uptake scenario with bidirectional charging, where generation from solar PV also decreases, and generation from bio energy increases (Figure 11). This is because the loss of bidirectional flexibility makes it favorable to use more dispatchable bio energy instead of a combination of more solar power and storage. Note that this switch from wind to solar power is not driven by the changes in BEV grid availability, as the latter is not the limiting factor even in the high uptake scenario (compare Figure 9).
@@
Figure 5: Yearly electricity generation by technology (“Shared + other BEVs”)
When private BEVs with smart charging are substituted with shared BEVs, we observe an increase in optimal long-duration storage and PV capacities, and lower firm conventional capacity with high variable costs (OCGT) (Figure 4). As the fleet with carsharing is more frequently used for driving, it offers less room to shift the demand to extended periods with high availability of RES (e.g., July 6-8 in Figure 16).
In the case where all BEVs charge bidirectionally, the switch to carsharing entails additional investments into firm conventional capacity, in particular OCGT (about 1 to 3 additional GW) and to a lesser extent also CCGT (about 0.5 GW). Yet, this does not translate into a greater yearly electricity generation from gas power plants (Figure 5). The additional firm capacity is only needed during some periods in which the residual load is particularly high, and which are served by V2G of private BEVs in the reference, but to a lower extent by shared BEVs. In contrast to the smart charging case, the switch to carsharing does not increase electricity storage needs when BEVs charge bidirectionally. Here, the remaining fleet of private BEVs still provides such a high level of flexibility that there is no need for any investments in short- or long-duration storage (compare also Figures 22 and  23).
This, however, is not the case of in the “Shared only setting”, where no additional private BEVs are available. Hence, we observe additional investments into long-duration storage discharging capacity of almost 3 GW for both uptake regimes as well as into short-duration storage of about 5.5 GW in the low uptake regime (Figures 20 and 21). The same holds for storage charging capacity (Figures 18 and 19) and storage energy capacity (Figures 22 and  23). For the latter, carsharing increases the optimal long-duration storage energy capacity by around 300 GWh for both uptake regimes. This closely resembles the overall battery capacity of substituted private BEVs (Table 2). In the case with smart charging, the effects of carsharing on optimal electricity storage capacities are much smaller. This is because vehicle batteries do not provide as much flexibility to the electricity sector in this setting, and the flexibility loss triggered by carsharing accordingly requires less compensation in terms of electricity storage.

### v2

@@ 2.3 Power sector modelling
We model four conventional generation technologies, including lignite, hard coal, open-cycle gas turbines (OCGT) and closed-cycle gas turbines (CCGT), as well as five renewable generation technologies, including solar photovoltaics, onshore wind, offshore wind, biomass and run-of-river hydropower (Table 7). Variable costs for conventional technologies include a carbon price of 130 euros per tonne of CO2eq. Three storage technologies are considered: lithium-ion batteries, pumped-hydro storage, and hydrogen-based long-duration storage; the latter being a combination of electrolysis, hydrogen storage, and hydrogen gas turbines (Table 8). A policy constraint reflects the renewable energy target of the German government: at least 80% of the electricity consumed has to come from renewable energy sources. Yet, this constraint is not binding in our parameterisation, as renewable shares around 85% lead to the lowest power sector costs (Table 10). Assuming plausible limits for expanding hydro and bio energy in Germany, the majority of this renewable electricity has to come from wind and solar power. In our parameterisation, this implies variable renewable energy shares of 80% or slightly more in all scenarios.
Modelling interactions with BEVs
Sector coupling with BEVs is modelled using distinct vehicle profiles. These profiles represent individual vehicles which are scaled up to the overall fleet. In the case of uncontrolled charging, each profile is considered an exogenous electricity demand that is added to the overall demand. When considering smart or bidirectional charging, the model chooses when and how much electricity should flow into a vehicle, while ensuring battery levels that are compatible with the mobility constraint, i.e. all trips until the next charging episode can be undertaken. This relies on the exogenous time series “driving electricity consumption” and “grid availability” derived with emobpy. Discharging electricity from BEV back to the grid incurs additional variable costs of 15 euros/MWh to account for battery degradation.
The model is calibrated based on a 15 million BEV fleet in 2030, which aligns with the target of the German coalition. The distribution of cars across locations is assumed to be the same as during the survey (Table 6). Additionally, it is assumed that clusters with very long trips (clusters 1 and 2 in each location with more than 250 kilometers driven on average during the day) are not regular daily patterns, but rather trips undertaken under exceptional circumstances, such as holidays. As these patterns represent a very small proportion of overall sequences (around 2.4%), we exclude them from our analysis and assume that the 15 million BEV fleet is distributed across locations and clusters 3 to 6 (Table 1). This means that very long trips that would constitute daily mobility habits are assumed not to be undertaken with electric cars by 2030. Nevertheless, very long trips can still take place sporadically during the day as well as over weekends.
@@ 2.4 Scenarios
We investigate a number of scenarios which differ with respect to the size of the overall vehicle fleet, the number of shared cars, and the charging strategy. Scenarios are compared to respective reference settings without carsharing, i.e. references only include privately owned BEV profiles.
Frameworks
We group our analyses into two broad frameworks. The first framework only considers BEVs that switch to carsharing in the scenario (“Shared only”). The second framework additionally includes other electric cars that stay private both in the reference and the scenario (“Shared + other BEVs”). Hence, the BEV fleets considered in the latter framework are larger than the ones in the former (Table 2). The former framework highlights the power sector effects of switching to carsharing, assuming away many other potential sources of flexibility. While results are instructive, they may not be policy-relevant. The latter framework is designed to mimic a more realistic setting with additional BEVs that do not switch to carsharing, but remain privately owned. Depending on the charging strategy, these additional “non-switching” BEVs also serve as a proxy for other flexibility options that are not included in the model. We thus consider “Shared + other BEVs” to be the main framework.
@@
Table 2: BEV units, battery capacity and power consumption for all reference settings and scenarios. Lecture note: the table displays the two frameworks defined above. For each framework, we consider several uptake regimes. For each framework and uptake regime, the reference consists in a BEV fleet without carsharing and the scenario consists in a fleet with carsharing. The overall number of BEVs in the fleet and its composition (private vs. shared BEVs) is detailed. The number of substituted BEVs corresponds to the number of BEVs that are removed from the fleet in the carsharing scenario compared to the reference. In the “Shared + other BEVs” framework with a high carsharing uptake, the reference fleet amounts to 15 million BEVs. In the scenario, 7.8 million private BEVs are substituted by 1.6 million shared BEVs. 7.2 million private BEVs remain in the fleet as private BEVs. The overall fleet is then composed of 8.8 BEVs in total, which corresponds to a removal of 6.2 million BEVs compared to the reference.
@@
For each framework, we consider two possible carsharing uptake regimes: a low one and a high one. In the low uptake regime, only BEVs belonging to the cluster of shortest trips (cluster 6) are assumed to switch to carsharing. In the high uptake regime, we assume that all BEVs in metropolises switch to carsharing as well as most BEVs in big cities and some in the middle-size cities. In small cities, only cluster 6 is assumed to switch to carsharing. For both uptake regimes, we assume that there is no switch to carsharing in rural areas. The number of substituted BEVs is approximately 3.9 million in the low uptake scenarios compared to 6.2 million in the high uptake scenarios. Further details are available in Figure 2 and Table 2. The low uptake regime is more conservative, as only users that rely the less on their cars switch to carsharing. The high uptake regime is more progressive and aims at illustrating a very optimistic scenario when analysing impacts on the power sector. The distribution of the number of daily trips in the overall fleet can be seen in Figure 9 and the distribution of where electric cars are located at each time step in Figure 10. Average grid availability for each hour of the week is illustrated in Figure 11.
Figure 2: Definition of carsharing uptake regimes based on cluster and location types. Note that this represents the reference fleet of the “Shared + other BEVs” framework. For the “Shared only” framework, the reference fleet is only constituted of the shaded areas that switch to shared BEVs in the scenario.
@@
For each scenario, we consider various charging strategies: uncontrolled charging, smart charging, or bidirectional charging. In order not to mix up effects, we assume that both the reference and the respective scenarios always have the same charging strategy. Additionally, only one charging strategy is used for the entire fleet at a time to enable meaningful comparisons. This means that all BEVs charge smartly in the reference and the scenarios, or they all charge bidirectionally, etc. We consider different charging strategies since it is unclear which charging strategy will be mostly used by BEVs in the future. While it is unlikely that the whole BEV fleet will be operated in an uncontrolled manner ([schill2015power]), it also seems unlikely that the whole fleet will be operated bidirectionally. By considering these strategies and homogeneous fleets, we analyse extreme settings which can be thought of as boundaries for outcomes of alternative scenarios with BEV fleets with mixed charging strategies.
@@ 3 Results
In the main part of the paper, we focus on the main framework where only part of the total BEV fleet considered in the reference switches to carsharing (“Shared + other BEVs”), i.e. the most policy-relevant framework. Results for the “Shared only” framework are provided in the appendix (D). Qualitative results are similar across frameworks unless explicitly mentioned.
@@ 3.1 Power sector cost increases remain moderate and are largest for bidirectional charging scenario
The adoption of carsharing in the BEV fleet results in increased power sector costs across almost all frameworks, carsharing uptake regimes, and charging strategies (Figure 3 and Table 9). This cost increase is relatively small compared to the overall sector costs, and always remains below 110 euros per substituted car per year in the main framework “Shared + other BEVs”, even in the case with bidirectional charging. In the “Shared + other BEVs” framework, it represents about 1.5% of the total sector costs and 2.2% in the “Shared only” framework. Note that the load of substituted BEVs is also small compared to the overall load and ranges between 1.1% and 2.6% of the overall load in the “Shared + other BEVs” framework, depending on the uptake regime. Sensitivity runs indicate that cost effects are qualitatively robust with respect to alternative weather years (Figures 30 and 31) and alternative technology cost assumptions (Figure 38 and Figure 39).
@@
Figure 3: Power sector costs by charging strategy and carsharing uptake [“Shared + other BEVs”]
@@
In the case where all BEVs charge in an uncontrolled way, switching to electric carsharing may even lead to slight cost reductions. This is because charging a large fleet of BEVs without considering power sector dynamics leads to a very expensive power sector, as BEVs considerably increase existing system peak loads (Figures 14 and 15). By switching to carsharing, the changing charging patterns may even slightly ease this burden. The cost decreases related to carsharing with the uncontrolled charging strategy should not be misinterpreted. The absolute sector costs remain the highest compared to other charging strategies, even after a switch to carsharing. Since the perspective of having a large BEV fleet charged in an uncontrolled way is neither realistic nor desirable (compare [schill2015power]), we do not focus on this charging strategy in the rest of the paper. Results can nevertheless be found in the appendix.
@@
Assuming that the whole BEV fleet charges smartly or bidirectionally, costs in the reference decrease compared to a setting with uncontrolled charging, as the former charging strategies help avoid costly load peaks (see section 3.2), which reduces the need for peaker plants and storage capacity. In turn, switching to carsharing then increases costs, regardless of the uptake regime. This is because substituting private BEVs that are charged smartly or bidirectionally with shared BEVs partly takes away the flexibility that private BEVs provide to the power sector, as shared cars are driven more often. Note that we assume that private cars provide power sector flexibility without additional investment costs. We also do not include charging infrastructure costs in our model, considering that the charging infrastructure assumptions are derived from exogenous mobility requirements. Substituting these private cars away implies an increase in the investment needed in the power sector. The cost increase is about two to three times higher when the whole BEV fleet is charged bidirectionally compared to when it is charged smartly, irrespective of the modelled uptake regime.
@@
For both the smart and the bidirectional charging strategy, the increase in cost is higher when the carsharing uptake is higher. In the “Shared + other BEVs” framework, this holds not only in absolute terms, but also in relative terms: the cost increase in euros per substituted car is always greater, everything else being equal, when considering the high uptake. It ranges between 35 and 49 euros per substituted BEV per year in the smart charging case, and between 99 and 108 euros for bidirectional charging
@@
When considering bidirectional charging, the increase in power sector costs related to carsharing is 40% to 75% higher in the “Shared only” framework, where no private BEV remain as a potential flexibility resource, compared to “Shared + other BEVs” (Table 9). This is because the remaining private BEVs (which are absent in the “Shared only” framework) are more likely to be used flexibly since they are driven less and are thus available for grid interactions more often. Feeding back electricity from BEV batteries into the grid is valuable in power sectors with high shares of renewables, as it helps to substitute peak generation and electricity storage capacities ([kern2022modeling, sovacool2017future]). Reducing the potential of a BEV fleet to engage in V2G accordingly hurts more than just reducing its potential to charge smartly. When assuming that all BEVs only charge smartly, but do not engage in V2G, the relative cost increases hardly differ between the “Shared only” and “Shared + other BEVs” frameworks.
@@
Figure 4: Generation and storage discharging capacity by technology [“Shared + other BEVs”]
@@ 3.2 Changes in generation capacity and dispatch depend on charging strategy
The switch to carsharing triggers changes in the optimal electricity generation capacity mix that are different for smart and bidirectional charging. In the case of smart charging, the switch to carsharing reduces the potential to flexibilize the demand. In the case of bidirectional charging, switching to carsharing also entails a loss in the ability of the BEV fleet to feed back electricity into the grid when the demand is relatively high compared to the available supply.
In the case of smart charging, the effects on the generation capacity generally remain small, with a moderate increase in onshore wind power investments of slightly less than 2 GW (Figure 4). This finding is qualitatively robust also for alternative weather years, where onshore wind capacity increases for most weather years by 1.7 GW on average, and only decreases in three of 22 weather years (Figure 33). Besides, we observe for all weather years except one additional PV investments but the magnitude of the additional capacity depends more strongly on the weather year, ranging from 0.4 to 7 GW with an average of 2.6 GW.
@@
For bidirectional charging, capacity effects are larger. The optimal PV generation capacity increases, by between around 1.5 and 5 GW. In contrast, optimal wind power capacities tend to decrease with carsharing (Figure 4, see also Figure 12). This finding holds for many of the alternative weather years (15 out of 21) and alternative cost assumptions (Figure 40 and Figure 41). This is because the relatively large cumulative battery capacity of the fleet of privately owned BEV tend to foster the integration of wind power in the reference. This is visible in the cumulative BEV battery charging level time series, which does not only show diurnal patterns related to PV variability, but is also dominated by longer-duration patterns that relate to wind power variability (Figures 26 and 27). If part of the BEV fleet switches to carsharing, the potential for such longer-duration storage decreases in line with the shrinking overall battery capacity. In turn, the share of wind power in the optimal capacity mix decreases, while the use of PV with its shorter-duration diurnal variations increases. Note that this switch from wind to solar power is not driven by the changes in BEV grid availability, as the latter is not the limiting factor even in the high uptake scenario (compare Figure 11).
@@
Figure 5: Yearly electricity generation by technology [“Shared + other BEVs”]
As the longer-duration storage capabilities of the car fleet decrease when private BEVs with bidirectional charging are substituted with shared BEVs, optimal investments into stationary long-duration electricity storage increase. This is true for both charging and discharging power (Figures 21 and 23) as well as storage energy capacity (Figure 25) of stationary long-duration storage. Depending on the weather year considered, we find substantial variations in the magnitude of the additional investment in long-duration storage. For the discharging capacity, the additional investment varies between 0.5 GW and 7.5 GW, with an average of 3 GW (Figure 33). For the energy capacity, the additional investment also shows important variations but is always positive, except for one year (Figures 36 and 37). Likewise, investments in firm CCGT and OCGT capacities increase (Figure 4). This generally holds for most weather year sensitivities (Figures 32 and 33). Yet, these additional capacities do not necessarily translate into greater yearly electricity generation from gas power plants (Figure 5 and Figures 34 and 35). In particular, generation from OCGT plants hardly increases, as some of the additional firm capacity is only needed during very few hours in which the residual load is particularly high, and which are served by V2G of private BEVs in the reference, but to a lower extent by shared BEVs. In contrast, electricity generation from CCGT plants increases (between one and four additional TWh for the year chosen in the main setting, see Figure 13 and Figures 42 and 43). This means that the renewable share slightly decreases and the electricity mix gets slightly more carbon-intensive. In the case of the “Shared only” setting, where no additional private BEVs are available for providing long-duration storage capabilities, we observe even higher investments into stationary storage charging and discharging capacity (Figures 20-23). In contrast to the bidirectional charging case, the switch to carsharing hardly impacts stationary electricity storage needs when BEVs only charge smartly. This is due to the fact that necessary storage requirements are already realized in the reference (Figures 20 to  25).